OpenStudy (anonymous):

Choose the correct simplification of 2a^0 ?

OpenStudy (turingtest):

\[a^0=1\text{ for all }a\neq0\]

OpenStudy (noseboy908):

a^0 =1, so 2x1=2

OpenStudy (anonymous):

omg idk which one cuz everyone gave me different answers

OpenStudy (cwtan):

For actual (2a)^0=1 2(a^0)=2

OpenStudy (anonymous):

okay guys thankyou soo muchhh

OpenStudy (anonymous):

Choose the correct simplification of